OpenStudy (amberparrishxo):

WILL MEDAL and FAN FOR CORRECT ANSWER! The prism has a volume of 120 cm3. What would the volume of the prism be if each dimension were halved?

OpenStudy (haseeb96):

V= a*c*h/2 since V=120 and if all the dimensions will be halved then we would write it as 120/6 = a*c*h/16 20cm3 is the correct answer

OpenStudy (haseeb96):

what do u think what is its answer?

OpenStudy (amberparrishxo):

At first i thought it was 15.. But this makes so much more since now.. Thank you!

OpenStudy (haseeb96):

Most welcome @amberparrishxo

OpenStudy (amberparrishxo):

For anyone coming later to get the correct answer, i took the test and got it incorrect.. The correct answer was infact 15 cm^3!